OpenStudy (anonymous):

A certain business has implemented a new incentive plan whereby an employee earns tokens for every sale they make. For 1 sale, an employee earns 1 token. For 2 sales, an employee earns 9 tokens. For 3 sales, an employee earns 21 tokens. For 4 sales, an employee earns 37 tokens. Which recursive equation represents the pattern?

OpenStudy (anonymous):

I think its an=an-1+(n-4) but not sure

OpenStudy (anonymous):

@jim_thompson5910 approval

OpenStudy (anonymous):

approved?*

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

Okiee :D

jimthompson5910 (jim_thompson5910):

well notice how the jump from 1 to 9 is +8 the jump from 9 to 21 is +12 the jump from 21 to 37 is +16

jimthompson5910 (jim_thompson5910):

so the jumps are: 8, 12, 16, ... and this pattern continues

OpenStudy (anonymous):

yes? but istn each one minus 4 backwards?

jimthompson5910 (jim_thompson5910):

so if the first term is a1 = 1 then you need to add 8 to it to get a2 = a1 + 8 a2 = 1+8 a2 = 9

jimthompson5910 (jim_thompson5910):

the third term is a3 = a2 + 12 a3 = 9 + 12 a3 = 21

jimthompson5910 (jim_thompson5910):

hopefully you're seeing a relationship form between the index number and the number we're adding on

jimthompson5910 (jim_thompson5910):

a4 = a3 + 16 a4 = 21 + 16 a4 = 37

OpenStudy (anonymous):

adding 4?

jimthompson5910 (jim_thompson5910):

it's not n-4 but here's a hint: 2*____ = 8 3*_____ = 12 4*_____ = 16 etc etc

OpenStudy (anonymous):

is it an=an-1+4n?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

rule: next term = (previous term) + 4*(index of next term)

OpenStudy (anonymous):

i have like 5 more questions, would you mind helping me out? :/

OpenStudy (anonymous):

Now i get it, i was just counting backwards, like an idiot

jimthompson5910 (jim_thompson5910):

that's quite a bit, but we'll see how things go

OpenStudy (anonymous):

Okay, thanks again @jim_thompson5910

jimthompson5910 (jim_thompson5910):

np